# FoREST: Frame of Reference Evaluation in Spatial Reasoning Tasks

### v2

@@ 4 Models and Tasks
This paper focuses on Question-Answering and Text-to-Image using the FoREST benchmark to evaluate FoR in spatial reasoning comprehensively. FoREST also supports other tasks, such as FoR identification, detailed in Appendix H.
@@ 4.1 Question-Answering (QA)
Task. This QA task evaluates LLMs’ ability to adapt contextual perspectives across different FoRs. The input includes a spatial expression $T$, relatum orientation (if available), and a question $Q$ querying the spatial relation from either the observer’s or the relatum’s perspective. The output is a spatial relation $S$, restricted to {left, right, front, back}. We use the context from both splits for this task.
@@
Chain-of-Thought (CoT) baseline (Wei et al., 2023). To create CoT examples, we modify the prompt to require reasoning before answering. We manually craft reasoning explanations with the necessary information for the few-shot examples. The input to the model is the instruction, CoT example, spatial context, and the question.
Human baseline. We conducted a small scope human study with three participants, paid as research assistants, to establish a human baseline for A-split. Each participant is shown 25 randomly sampled QA examples per relatum case, totaling 150 examples. Each includes a textual scene description and corresponding question. See Appendix I for details. We report average accuracy across participants to reflect the human baseline.
Camera perspective Relatum perspective
@@
Table 1: QA accuracy in the A-Split. Rand Irepresent the percentage the model assumes relative or intrinsic FoR for ambiguous expression, explained in Section 5.1. Acc is the accuracy, and Avg is the micro-average of accuracy. (1): 0-shot, (2): 4-shot, (3): CoT, and (4): SG+CoT.
@@ 4.2 Text-To-Image (T2I)
Task. This task assesses the diffusion models’ ability to consider FoR by evaluating their generated images. The input is a spatial expression, $T$, and the output is a generated image ($I$). C and A splits with external FoRs are used for this task.
Stable Diffusion Models. We use the stable diffusion models as the baseline for the T2I task. This model only needs the scene description as input.
Layout Diffusion Models. This model operates in two phases: text-to-layout and layout-to-image. Given that LLMs can generate the bounding box layout (Cho et al., 2023b), we provide them with instructions and $T$ to create the layout. The layout consists of bounding box coordinates for each object in the format of {object: $[x,y,w,h]$}, where $x$ and $y$ denote the starting point and $h$ and $w$ denote the height and width. The bounding box coordinates and $T$ are then passed to the layout-to-image model to produce the final image, $I$.
@@ 4.3 Spatial-Guide (SG) Prompting
We hypothesize that the spatial relation types and FoR classes defined in Section 2 can improve QA and layout generation. For example, the external intrinsic FoR emphasizes that spatial relations originate from the relatum’s perspective. To leverage this, we propose SG prompting, an additional step applied before QA or layout generation. This step extracts spatial information, including direction, topology, distance, and the FoR, from the spatial expression $T$, as supplementary input to guide LLMs in QA or layout generation. We manually craft four examples covering these aspects. First, we specify the perspective for directional relations, e.g., left relative to the observer, to distinguish intrinsic from relative FoR. Next, we indicate whether the locatum is inside or outside the relatum for topological relations to differentiate internal from external FoR. Lastly, we provide an estimated quantitative distance to support topological and directional relation identification, e.g., far. These examples are provided in a few-shot setting to guide the model in automatically extracting such information.
@@ 5 Experimental Results
Camera perspective Relatum perspective
@@
Table 2: QA accuracy in the C-Split across various LLMs. ER, EI, II, and IR denote external relative, external intrinsic, internal intrinsic, and internal relative FoRs. Avg represents the micro-average accuracy. CP refers to context with camera perspective, while RP denotes context with relatum perspective.
@@
Llama3-70B + SG + GLIGEN (Our) $56.54$ $30.59$ $\mathbf{87.13}$ $66.56$ $\mathbf{83.75}$ $56.77$ 70.0470.04{70.04}70.04 64.0664.06{64.06}64.06
Table 3: VISORcond score explained in Section 5.1 where $I$ refers to the Cow and Car cases where relatum has intrinsic directions, and $R$ refers to the Box and Pen cases where relatum lacks intrinsic directions, $avg$ is the micro-average of $I$ and $R$. EI and ER FoR represent the generated image considered corrected by EI or ER FoR.
@@ 5.1 Evaluation Metrics
QA. We report an accuracy (acc.) based on the correct answer defined as follows. Since the questions can have multiple correct answers, as explained in Section 3, the prediction is correct if it matches any valid answer. Additionally, we report the model’s bias distribution when FoR ambiguity exists. $I$is the percentage of correct answers when assuming an intrinsic FoR, while $R$is this percentage with a relative FoR assumption. Note that cases where both FoR assumptions lead to the same answer are excluded from the bias calculation.
@@ 5.3 Results
RQ1. What is the LLM’s bias for the ambiguous FoR? Table 1 presents the QA results for the A-split. Ideally, a model extracting spatial relations without considering perspective should achieve 100% accuracy, as the context lacks a fixed perspective. However, our focus is not on this. We aim to assess LLMs’ bias by measuring how often they adopt a specific perspective when answering. In the Box and Pen cases, relatum properties do not introduce FoR ambiguity in directional relations, making the task pure extraction rather than reasoning. Thus, we focus on the $I$and $R$of the Cow and Car cases, which best reflect LLMs’ bias. Qwen2 achieves 80% accuracy across all experiments by selecting spatial relations directly from context, suggesting it may disregard the question’s perspective. This is supported by attention analysis in Appendix C. The attention maps show that most correct Qwen2 responses assume a shared perspective between context and question and pay low attention to the perspective mentioned in the question. GPT-4o exhibits similar bias in 0-shot and 4-shot settings but shifts toward intrinsic interpretation with CoT. This shift reduces accuracy in camera-perspective questions where FoR adaptation is more significant than relation extraction. Llama3-70B shows no strong preference, slightly favoring relative FoR, but this balanced assumption leads to lower performance due to increased reasoning demands. Qwen2VL exhibits a similar pattern but achieves higher accuracy, likely due to visual training. GPT-o4-mini with high visual reasoning shows the same pattern, which is later discussed as an additional experiment in Appendix J. Moreover, our experiments with humans, presumed to have prior visual knowledge, show strong reasoning ability, achieving over 90% accuracy even in scenarios where perspective preference requires FoR adaptation. In our results, humans slightly prefer intrinsic FoR over relative FoR. Their cultural background and recently seen examples may also influence their preference as discussed in Coventry et al. (2018). A more detailed analysis of human performance is provided in Appendix I. In summary, we conjecture that Qwen2 performs well by focusing on extraction without reasoning FoR, while other models attempt reasoning but often fail to reach correct conclusions, leading to lower accuracy.
@@
Figure 3: Confusion matrices of predicted spatial relations by Llama3 and GPT-4o when adapting to FoR is required. 0-shot and SG-COT settings are reported here.
@@
RQ2. Can LLMs adapt FoR when answering the questions? To address this RQ, we analyze QA results of C-split in Table 2, where context and question explicitly specify perspective. Results show that LLMs struggle with FoR conversion, achieving only up to 55.5% accuracy when the question uses the relatum’s perspective and the context uses the camera’s. We further examine how Llama3 and GPT-4o adapt FoR using the confusion matrix in Figure 3. Our findings reveal that pure-text LLM (Llama3) reverses left and right. This contrasts with humans when describing the spatial relation from a perspective in English, which reverses front and back while preserving lateral directions. This leads to poor adaptation to the camera perspective of Llama3. In contrast, large multimodal models like GPT-4o follow expected patterns, consistent with Zhang et al. 2025. While our GPT-4o results suggest some abilities to convert the relatum’s perspective into the camera’s with in-context learning (72% accuracy), the reverse transformation in the textual domain remains challenging (53% accuracy). A similar trend appears when comparing Qwen2 and Qwen2VL, as discussed in Appendix B. Qwen2VL performs better when shifting to the camera perspective but performs worse with reverse transformation, often failing to generate correct reasoning. This difficulty persists for several models when converting spatial relations from images to relatum’s perspective, as noted in Zhang et al. 2025. A possible reason is that the model often overlooks orientation tokens, critical for answering FoR questions, based on Attention analysis provided in Appendix C.
@@
RQ3. How can an explicit FoR identification help spatial reasoning in QA? We compare CoT and CoT+SG results to assess the effect of FoR identification on LLMs’ spatial reasoning in QA. Based on C-Split results (Table 2), incorporating SG improves the model’s ability to identify the correct perspective from input expression ranging from 2.9% to 30% of cases where the context and question share the same perspective. These cases are easier as the models do not need FoR adaptation. Two notable exceptions emerge. First, Llama3 performs poorly on camera-perspective questions, and FoR identification via SG fails to improve its performance. This may be due to Llama3’s lack of visual training, which we speculate limits its FoR understanding. Second, Qwen2VL struggles with relatum-perspective reasoning, showing negative gains even with CoT. SG is less effective when context and question differ in perspective; while it helps identify the correct FoR in context, it does not enhance reasoning across perspectives. This limitation is evident in A-Split results (Table 1), where models only improve significantly when SG aligns their preference with the question’s perspective, as seen in Qwen2-72B and GPT-4o. SG identification results are reported in the Appendix H. Still, FoR identification improves overall spatial reasoning (see Avg. column for SG+CoT in Table 3).
@@
RQ4. How can explicit FoR identification help spatial reasoning in visualization? We evaluate SG layout diffusion to assess the impact of incorporating FoR in image generation. We focus on VISORcond metric, which better reflects the model’s spatial understanding than the overall performance measured by VISORuncond, reported in Appendix F. Table 3 shows that adding spatial information and FoR classes (SG+GLIGEN) improves performance across all splits compared to the baseline models (GLIGEN). In particular, SG improved the model’s performance when expressions follow a relative FoR. This finding aligns with QA results in Table 1, where Llama3 prefers relative FoR in camera-perspective scenarios. In contrast, baseline diffusion models (SD-1.5 and SD-2.1) perform better for intrinsic FoR, even though GLIGEN is based on SD-2.1. This outcome may be due to GLIGEN’s reliance on bounding boxes for spatial configurations, which makes it struggle with intrinsic FoR due to the absence of object properties and orientation. Despite this bias, incorporating FoR information via SG-prompting improves performance across all FoR classes. We provide further analysis on SG for the layout generation in Appendix G.
@@ 6 Related Work
Frame of Reference in Cognitive Studies. The concept of the frame of reference in cognitive studies was introduced by Levinson 2003 and later expanded with more diverse spatial relations (Tenbrink, 2011). Subsequent research investigated the human preferences for specific FoR classes (Vukovic and Williams, 2015; Shusterman and Li, 2016; Ruotolo et al., 2016; Coventry et al., 2018). For instance, Ruotolo et al. 2016 examined how FoR influences scene memorization. Their study found that participants performed better when spatial relations were based on their position rather than external objects, highlighting a distinction between relative and intrinsic FoR.
Frame of Reference in AI. Several benchmarks have been developed to evaluate the spatial understanding of AI models in multiple modalities; for instance, textual QA (Shi et al., 2022; Mirzaee and Kordjamshidi, 2022; Rizvi et al., 2024), and text-to-image (T2I) benchmarks (Gokhale et al., 2023; Cho et al., 2023a, b). However, most of these benchmarks overlook FoR, assuming a single FoR for all instances despite its significance in cognitive studies. Recent vision-language studies have begun addressing this gap Liu et al. (2023); Zhang et al. (2025). For instance, Zhang et al. 2025 explores FoR ambiguity by evaluating spatial relations from camera-perspective images, with FoR specified in the question. In contrast, our work focuses on the spatial reasoning when dealing with multiple FoRs and when there are changes in perspective in explaining the context beyond the camera’s viewpoint. We further show that explicitly identifying FoR enhances spatial reasoning in QA and T2I.

### v3

@@ 4 Models and Tasks
This paper focuses on Question-Answering and Text-to-Image tasks using the FoREST benchmark to evaluate FoR in spatial reasoning comprehensively. FoREST supports additional tasks, such as FoR identification, detailed in Appendix K.
@@ 4.1 Question-Answering (QA)
Task. The QA task evaluates LLMs’ ability to adapt contextual perspectives across different FoRs. The input includes a spatial expression $T$, relatum orientation (if available), and a question $Q$ querying the spatial relation from either the observer’s or the relatum’s perspective. The output is a spatial relation $S$, restricted to {left, right, front, back}. We use both dataset splits for this task.
@@
Chain-of-Thought (CoT) baseline (Wei et al., 2022). To create CoT examples, we modify the prompt to require reasoning before answering. We manually craft reasoning explanations with the necessary information for the few-shot examples. The input to the model is the instruction, CoT example, spatial context, and the question.
Human baseline. We conducted a small-scale human study with three participants, compensated as research assistants. Each participant was shown 25 randomly selected QA examples per relatum case in the A-split, for a total of 150 examples. Each example consisted of a textual scene description and a corresponding question. See Appendix G for details. We report average accuracy across participants to reflect the human baseline.
Question with Camera Perspective Question with Relatum Perspective
@@
Table 2: QA accuracy in the A-split. R% and I% indicate the proportion of cases where the model assumes a relative or intrinsic FoR for an ambiguous expression (see Section 5.1). Acc denotes accuracy, and Avg is the micro-average accuracy. (1) 0-shot, (2) 4-shot, (3) CoT, and (4) SG+CoT.
@@ 4.2 Text-To-Image (T2I)
Task. This task assesses the ability of diffusion models to consider FoR by evaluating their generated images. The input is a spatial expression, $T$, and the output is a generated image ($I$). C and A splits with external FoRs are used for this task.
Stable Diffusion Model. We use the stable diffusion model as the baseline for the T2I task. This model only needs the scene description as input.
Layout Diffusion Model. This model operates in two phases: text-to-layout and layout-to-image. Given that the LLM can generate the bounding box layout (Cho et al., 2023b), we provide the LLM with instructions and $T$ to create the layout. The layout consists of bounding box coordinates for each object in the format of {object: [x,y,w,h][x,y,w,h]}, where $x$ and $y$ denote the starting point and $h$ and $w$ denote the height and width. The bounding box coordinates and $T$ are then passed to the layout-to-image model to produce the final image, $I$.
@@ 4.3 Spatial-Guide (SG) Prompting
We hypothesize that the spatial relation types and FoR classes defined in Section 2 can guide QA and layout generation. For example, the external intrinsic FoR emphasizes that spatial relations originate from the relatum’s perspective. To leverage this, we propose SG prompting, an additional step applied before QA or layout generation. This step extracts spatial information, including direction, topology, distance, and the FoR, from the spatial expression $T$, as supplementary input to guide LLMs in QA or layout generation. We manually craft four examples covering these aspects. First, we specify the perspective for directional relations, e.g., left relative to the observer, to distinguish intrinsic from relative FoR. Next, we indicate whether the locatum is inside or outside the relatum for topological relations to differentiate internal from external FoR. Lastly, we provide an estimated quantitative distance to support topological and directional relation identification, e.g., far. These examples are provided in a few-shot setting to guide the model in automatically extracting such information. The extracted information is then used to guide CoT reasoning in QA and layout generation.
@@ 5 Experimental Results
Question with Camera Perspective Question with Relatum Perspective
@@
Table 3: QA accuracy in the C-Split across various LLMs. ER, EI, II, and IR denote external relative, external intrinsic, internal intrinsic, and internal relative FoRs, respectively. Avg denotes the micro-average accuracy. CP indicates context with a camera perspective, while RP denotes context with a relatum perspective.
@@
Llama3-70B + SG + GLIGEN (Our) $56.54$ $30.59$ $\mathbf{87.13}$ $66.56$ $\mathbf{83.75}$ $56.77$ 70.04{70.04} 64.06{64.06}
Table 4: VISORcond score explained in Section 5.1 where II refers to the Cow and Car cases where relatum has intrinsic directions, and $R$ refers to the Box and Pen cases where relatum lacks intrinsic directions, avgavg is the micro-average of II and $R$. EI and ER FoR represent the generated image considered corrected by EI or ER FoR.
@@ 5.1 Evaluation Metrics
QA. We report accuracy (acc.) based on the correct answer defined as follows. Since the questions can have multiple correct answers, as explained in Section 3, the prediction is correct if it matches any valid answer. Additionally, we report the model’s bias distribution when FoR ambiguity exists. II% is the percentage of correct answers when assuming an intrinsic FoR, while RR% is this percentage with a relative FoR assumption. Note that cases where both FoR assumptions lead to the same answer are excluded from the bias calculation.
@@ 5.3 Results
RQ1. What is the LLM’s bias when FoR is ambiguous? Table 2 presents the QA results for the A-split. Since the context lacks a fixed perspective, a model extracting spatial relations alone should ideally reach 100% accuracy. Our goal, however, is to assess LLMs’ bias by measuring how often they adopt a specific perspective. In the Box and Pen cases, relatum properties do not introduce FoR ambiguity in directional relations, making the task pure extraction rather than reasoning. Thus, we focus on the II% and RR% of the Cow and Car cases, which best reflect LLMs’ bias. Qwen2-72B consistently achieves 80–95% accuracy across all experiments by selecting spatial relations directly from context, suggesting it may disregard the question’s perspective. This is supported by attention analysis in Appendix D. The attention maps show that most correct Qwen2 responses assume a shared perspective between context and question and pay low attention to the perspective stated in the question. GPT-4o exhibits similar bias in 0-shot and 4-shot settings but shifts toward intrinsic interpretation with CoT. This shift reduces accuracy on camera-perspective questions, where FoR adaptation plays a larger role than relation extraction. Llama3-70B shows no strong preference, slightly favoring relative FoR, but this balance lowers performance due to increased reasoning demands. Qwen2VL follows a similar pattern but achieves higher accuracy, likely due to visual training. GPT-o4-mini with high visual reasoning shows the same pattern, which is later discussed as an additional experiment in Appendix C. Moreover, our experiments with humans, presumed to have prior visual knowledge, show strong reasoning ability, achieving over 90% accuracy even in scenarios where perspective preference requires FoR adaptation. In our results, humans slightly prefer intrinsic FoR over relative FoR. Their cultural background and recently seen examples may also influence their preference as discussed in Coventry et al. (2018). A detailed analysis of human performance is provided in Appendix G. In summary, we conjecture that Qwen2 performs well by focusing on extraction without reasoning FoR, while other models attempt reasoning but often fail to reach correct conclusions, leading to lower accuracy. To support this claim and further analyze the results, we provide a quantitative analysis in Appendix E, examining how linguistic expressions of spatial relations and facing directions in the context influence model outputs.
@@
Figure 3: Confusion matrices of spatial relation predictions by Llama3 and GPT-4o in 0-shot and SG+CoT settings, when FoR adaptation is required.
@@
RQ2. Can LLMs adapt FoR when answering the questions? To address this question, we analyze QA results of C-split in Table 3, where context and question explicitly specify the perspective. Results show that LLMs struggle with FoR conversion, especially when the question is asked from the relatum’s perspective and the context is given from the camera’s perspective, with the highest accuracy only 55.5% with Llama3-70B (CoT). We further analyze FoR adaptation in Llama3-70B and GPT-4o using the confusion matrix in Figure 3. Our findings reveal that pure-text LLM, Llama3-70B, systematically reverses left and right. This contrasts with humans, who in English typically reverse front and back when describing spatial relations from a perspective, while preserving lateral directions Hill (1982). This difference explains Llama3’s poor adaptation to the camera perspective. In contrast, large multimodal models like GPT-4o follow expected patterns, consistent with Zhang et al. 2025b. While the GPT-4o results suggest some ability to convert the relatum’s perspective into the camera’s with in-context learning (up to 82.4% accuracy), the reverse direction—from the camera’s perspective to the relatum’s—remains challenging, reaching only 53% accuracy (see GPT-4o+CoT). A similar trend appears when comparing Qwen2 and Qwen2VL, as discussed in Appendix B. Qwen2VL performs better when shifting to the camera perspective but performs worse with the reverse direction, often failing to generate correct reasoning. This difficulty persists for several models when converting spatial relations from images to the relatum’s perspective, as noted in Zhang et al. 2025b. Based on our Attention analysis, one possible reason for poor performance in perspective shift in Qwen2 is that the model often overlooks orientation tokens, critical for answering FoR questions. The detail of the analysis and visualizations can be found in Appendix D. Another observed reason is that the model tends to rely on a fixed facing direction when answering (facing toward or facing away). We provide a quantitative analysis of this behavior in Appendix E.
@@
RQ3. How can an explicit FoR identification help spatial reasoning in QA? We compare CoT and SG+CoT results to assess the effect of FoR identification on LLMs’ spatial reasoning in QA. Based on C-Split results (Table 3), incorporating SG improves the model’s ability to identify the correct perspective from input expression ranging from 2.9% to 30% of cases where the context and question share the same perspective. These cases are easier as the models do not need FoR adaptation. However, two notable exceptions arise. First, Llama3 performs poorly on camera-perspective questions, and FoR identification via SG fails to improve its performance. This may be due to Llama3’s lack of visual training, which we speculate limits its FoR understanding. Second, Qwen2VL struggles with relatum-perspective reasoning, showing negative gains even with CoT. SG is less effective when context and question differ in perspective; while it helps identify the correct FoR in context, it does not enhance reasoning across perspectives. This limitation is evident in A-Split results (Table 2), where models only improve significantly when SG aligns their preference with the question’s perspective, as seen in Qwen2-72B and GPT-4o. SG identification results are reported in the Appendix K. Still, FoR identification improves overall spatial reasoning (see Avg. column for SG+CoT in Table 4).
@@
RQ4. How can explicit FoR identification help spatial reasoning in visualization? To address this question, we evaluate the SG+GLIGEN baseline with a focus on the VISORcond metric, which better reflects the model’s spatial understanding than the overall performance measured by VISORuncond alternative that is reported in Appendix I. Table 4 shows that adding spatial information and FoR classes (SG+GLIGEN) improves performance across all splits compared to the GLIGEN baseline. In particular, SG improved the model’s performance when expressions follow a relative FoR. This finding aligns with QA results in Table 2, where Llama3 prefers relative FoR in camera-perspective scenarios. In contrast, baseline diffusion models (SD-1.5 and SD-2.1) perform better for intrinsic FoR, even though GLIGEN is based on SD-2.1. This outcome is likely due to GLIGEN’s reliance on bounding boxes for spatial configurations, which limits handling of intrinsic FoR because object properties and orientation—critical for intrinsic reasoning—are missing. Despite this limitation, incorporating FoR information via SG-prompting improves performance across all FoR classes. We provide further analysis on SG for the layout generation in Appendix J.
RQ5. How does human FoR understanding compare to LLMs’? For understanding FoR, humans usually adopt the perspective of a specific object, a process linked to the theory of mind, when describing spatial relations Loy and Demberg (2023). However, the ability to shift spatial perspective can be influenced by factors such as social context Creem-Regehr et al. (2013), situational circumstances Gunalp et al. (2019), and linguistic or cultural background Bohnemeyer et al. (2014). These factors highlight the inherent difficulty of FoR reasoning, even for humans. For LLMs, the challenge is even more pronounced. A key limitation stems from the nature of their training data, which is often based on image–caption pairs from 2D images or spatial descriptions grounded primarily in the visual modality from a camera perspective. As a result, models tend to learn spatial relations only from this specific viewpoint, restricting their ability to generalize beyond it. As shown in Table 2, our results indicate that most language-only models (e.g., Llama3) struggle with perspective shifts, whereas multimodal LLMs (e.g., GPT-4o) perform significantly better, though primarily from the camera perspective. This bias of MLLMs toward a single FoR system is consistent with previous findings Zhang et al. (2025b), which show that MLLMs often fail to adapt to cultural variation in perspective-taking, typically aligning with the English language. This bias underscores the need to improve LLMs’ ability to generalize across diverse spatial reasoning tasks. Future work would be interesting to explore novel ideas for both large-scale and realistic training data, as well as reasoning techniques that enable models’ FoR reasoning.
@@ 6 Related Work
Frame of Reference in Cognitive-Linguistic (CL) Studies. The concept of the frame of reference in CL studies was introduced by Levinson 2003 and later expanded with more diverse spatial relations (Tenbrink, 2011). Subsequent research investigated the human preferences for specific FoR classes (Vukovic and Williams, 2015; Shusterman and Li, 2016; Ruotolo et al., 2016; Coventry et al., 2018). For instance, Ruotolo et al. 2016 examined how FoR influences scene memorization. They found that participants performed better when spatial relations were based on their own position rather than external objects, highlighting a distinction between relative and intrinsic FoR.
Frame of Reference in AI. Several benchmarks have been developed to evaluate the spatial understanding of AI models in multiple modalities; for instance, textual QA (Shi et al., 2022; Mirzaee and Kordjamshidi, 2022; Rizvi et al., 2024), and text-to-image (T2I) benchmarks (Gokhale et al., 2023; Cho et al., 2023a, b). However, most of these benchmarks overlook FoR, assuming a single FoR for all instances despite its significance in cognition. Recent vision-language studies have begun addressing this gap Liu et al. (2023); Zhang et al. (2025b, a); Wang et al. (2025). For instance,Zhang et al. 2025a; Wang et al. 2025 propose benchmarks that incorporate the concept of perspective in embodied AI and situated 3D environments, although FoR is not the primary focus of either work. Liu et al. 2023 examines FoR’s impact on visual question answering but focuses only on the intrinsic and relative FoR categories. Our work covers a wider range of FoRs. Zhang et al. 2025b explores FoR ambiguity by evaluating spatial relations from camera-perspective images, with FoR specified in the question. In contrast, our work examines spatial reasoning across multiple FoRs and perspective changes, extending beyond the camera’s viewpoint. We further demonstrate that explicitly identifying FoR for in-context learning enhances spatial reasoning in both QA and T2I tasks.